OpenStudy (anonymous):

I realy need help ill give a madal. What is the volume of the cylinder? Use three point one four for pi. A cylinder with a diameter of 6 miles and a height of ten miles one hundred eighty-eight point four cubic miles two hundred forty-four point nine two cubic miles two hundred eighty-two point six cubic miles four hundred seventy-one cubic miles

OpenStudy (sidsiddhartha):

V = (pi)(r^2)(h) r = d/2 = 3 V = (3.14)(3^2)(10) = 282.6mi^3

OpenStudy (anonymous):

The volume of a cylinder has the equation\[V=\pi r^2h\]just use pi = 3.14 and plug in the heights you are given for h, and the radii for r

OpenStudy (anonymous):

\[3.14 * 3^2 * 10\] \[3.14 * 9 * 10\] \[3.14 * 90\] \[282.6\]

OpenStudy (anonymous):

Thanks so much guys ill give you a medal.